# On the asymptotic behavior of the spectral gap for discrete Schrödinger operators

### v1

@@ 3. Spectral estimates
$\psi_{k}(j):=\left(1-b_{k}\right)^{1/2}\left(\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\right)+\widehat{\alpha}^{-1/2}b^{1/2}_{k}c_{k}\ ,\quad j\in V_{k}\ ,\quad k\in\mathds{N}\ ,$
is normalized. Furthermore, there exists a constant $c_{1}>0$ independent of $k$ and $\boldsymbol{\alpha}$ such that
$b_{k}\geq c_{1}\widehat{\alpha}^{-1}k^{-1}$ (14)
@@
$\psi_{k}(j)=\left(1-b_{k}\right)^{1/2}\left(\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\right)+b^{1/2}_{k}\widehat{c}_{k}\ ,\quad j\in V_{k}\ ,\quad k\in\mathds{N}\ .$
@@
$b_{k}$ $=2\left(1-b_{k}\right)^{1/2}b_{k}^{1/2}\widehat{c}_{k}\sum\limits_{j=-k}^{k}\left(\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\right)+(2k+1)b_{k}\widehat{c}_{k}^{2}\ .$
@@
Comparing the upper bound for $\lambda_{0}(V_{k},\boldsymbol{\alpha})$ with the lower bound in Theorem 2, we conclude that, for all but finitely many $k\in\mathds{N}$, $b_{k}\leq(const.)(a_{k,1}+a_{k_{2}})$ and therefore $\lim_{k\to\infty}b_{k}=0$ by Theorem 2.
Consequently, for all but finitely many $k\in\mathds{N}$,
$b_{k}$ $\geq b_{k}^{1/2}\widehat{c}_{k}\sum\limits_{j=-k}^{k}\big{(}\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\big{)}\ .$
Hence, using that $\sum\limits_{j=-k}^{k}\big{(}\widetilde{\cos}_{k,r_{\min}}(j)+\widetilde{\cos}_{k,r_{\max}}(j)\big{)}\geq(const.)k^{1/2}$ for all but finitely many $k\in\mathds{N}$ and (5), we obtain
$b_{k}\geq(const.)\widehat{\alpha}^{-1}k^{-1}\ .$
@@
Lemma 6 (Bounds on the first excited state energy).
For all but finitely many $k\in\mathds{N}$ one has
@@
The lower bound is a direct consequence of $\lambda_{1}(V_{k},\boldsymbol{\alpha})\geq\lambda_{1}(V_{k},0)=\lambda_{0}(V_{k},\infty\delta_{0})$. The upper bound follows from the operator inequality $\mathcal{L}_{k}+\sum_{j\in J}\alpha_{j}\delta_{j}\leq\mathcal{L}_{k}+\sum_{j\in J}\infty\delta_{j}$ in combination with $\lambda_{1}(V_{k},\sum_{j\in J}\infty\delta_{j})\leq\max\{\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\}$ that holds for all but finitely many $k\in\mathds{N}$. ∎
4. Main results
@@
$\left(|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\right)_{k\in\mathds{N}}$
and prove a conjecture regarding its limiting behavior as put forward in [KY25]. By doing this and by considering potentials of compact support, we also generalize the main result of [KY25] – both – in terms of convergence speed and applicability to a broader class of potentials: more explicitly, we will show that
@@
for all $\eta<1$, as soon as the external potential is not the zero potential. Here, one should recall that $\lim_{k\rightarrow\infty}\left(|V_{k}|^{2}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\right)=\pi^{2}$ for $\boldsymbol{\alpha}=0$; compare with [KY25]. In other words, as soon as there is a non-vanishing external potential of compact support, the spectral gap converges strictly faster to zero than in the absence of any potential. This might be surprising at first, since the potential is supported on a smaller and smaller fraction of the configuration space in the limit of infinite volume.
We now state the main result of this paper.
@@
    There exists a constant $c_{1}=c_{1}(\boldsymbol{\alpha})>0$ such that
    $|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\leq c_{1}\quad\text{for all}\quad k\in\mathds{N}\ .$
@@
    If $\widehat{\alpha}=\sum_{j\in J}\alpha_{j}$ is sufficiently small or if $J=\{0\}$, there exists a constant $c_{2}=c_{2}(\boldsymbol{\alpha})>0$ such that
    $|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\geq c_{2}\quad\text{for all}\quad k\in\mathds{N}\ .$
@@
$|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\leq(const.)\widetilde{\alpha}^{-1}k^{-1}\Theta_{\max}|V_{k}|^{3}+\left(\Theta_{\max}-\Theta_{\min}\right)|V_{k}|^{3}\ .$ (15)
Now, taking into account that $\Theta_{\max}\sim k^{-2}$ and $\Theta_{\max}-\Theta_{\min}\lesssim k^{-3}$ completes the proof; note here that $\lambda_{0}(V_{m},\infty\delta_{0})=2-2\cos(\pi/|V_{m}|)$, $m\in\mathds{N}$; see, for example, [KY25].
@@
$|V_{k}|^{3}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\geq(const.)\widehat{\alpha}^{-1}\lambda_{0}(V_{k},\infty\delta_{0})k^{-1}|V_{k}|^{3}+\left(\lambda_{0}(V_{k},\infty\delta_{0})-\Theta_{\max}\right)|V_{k}|^{3}\ .$ (16)
@@
    Regarding the upper bound (15): Note that the first term is of the form $(const.)\widetilde{\alpha}^{-1}$ and consequently converges to zero when the potential strength converges to infinity in the sense that $\widetilde{\alpha}\to\infty$. However, the last term is independent of the potential strength and is non-zero if and only if $r_{\max}\neq-r_{\max}$. Therefore, the entire upper bound is of the form $(const.)\widetilde{\alpha}^{-1}$ and consequently converges to zero as $\widetilde{\alpha}$ converges to infinity if and only if $r_{\max}=-r_{\min}$.
@@
    Regarding the lower bound (16): The last term is negative whenever $J\neq\{0\}$ and it vanishes if and only if $J=\{0\}$; note that, trivially, $\Gamma(V_{k},\boldsymbol{\alpha})\geq 0$ for all $k\in\mathds{N}$. Therefore, this lower bound is non-trivial if and only if $\widehat{\alpha}$ is sufficiently small or if $J\in\{0\}$. Furthermore, the first term of this bound is of the form $(const.)\widehat{\alpha}^{-1}$. Consequently, if and only if $J=\{0\}$, the entire lower bound is of the form $(const.)\widehat{\alpha}^{-1}$ and converges to zero as the strength of the potential tends to infinity in the sense that $\widehat{\alpha}\to\infty$. Note that the lower bound converges to $\infty$ as $\hat{\alpha}\to 0$. This refers to the case where we approach the free Laplacian, for which the spectral gap vanishes in lower order.
@@
    The reason we end up with a trivial lower bound when $J\neq\{0\}$ and $\widehat{\alpha}$ is too large is the lack of a sufficiently good lower bound for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ in terms of $\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})$ and/or $\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})$; compare with Lemma 6. Indeed, also from the physical point of view outlined in Remark 4, it is reasonable to expect that the lower bound for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ in Lemma 6 is suboptimal for strong external potentials. On the other hand, however, a corresponding statement as in $(ii)$ of Theorem 7 is expected for all potential strengths.
We conclude this paper with the following corollary which summarizes our findings for the special case $J=\{0\}$ addressed in [KY25], and which originally motivated this work. It also proves the conjecture put forward in [KY25, Section 4].
@@
Consider the Hamiltonian $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with any $\alpha>0$ and associated spectral gap $\Gamma(V_{k},\alpha)$. Then there exist constants $c_{1},c_{2}>0$ independent of $k$ and $\alpha$ such that for all but finitely many $k\in\mathds{N}$,
$\dfrac{c_{1}}{\alpha}\leq|V_{k}|^{3}\cdot\Gamma(V_{k},\alpha)\leq\dfrac{c_{2}}{\alpha}\ .$
@@
$|V_{k}|^{3}\cdot\Gamma(V_{k},\alpha)\leq(const.)\alpha^{-1}k^{-1}\lambda_{0}(V_{k},\infty\delta_{0})|V_{k}|^{3}\ .$
@@
$|V_{k}|^{3}\cdot\Gamma(V_{k},\alpha)\geq(const.)\alpha^{-1}k^{-1}\lambda_{0}(V_{k},\infty\delta_{0})|V_{k}|^{3}$

### v2

@@ 3. Spectral estimates
$\psi_{k}(j):=\left(1-b_{k}\right)^{1/2}\left(\widetilde{\cos}_{k,r_{\min}}^{\text{l}}(j)+\widetilde{\cos}_{k,r_{\max}}^{\text{r}}(j)\right)+\widehat{\alpha}^{-1/2}b^{1/2}_{k}c_{k}\ ,\quad j\in V_{k}\ ,\quad k\in\mathds{N}\ ,$
is normalized. Furthermore,
$b_{k}\geq(const.)\widehat{\alpha}^{-1}k^{-1}$ (19)
@@
$\psi_{k}(j)=\left(1-b_{k}\right)^{1/2}\left(\widetilde{\cos}_{k,r_{\min}}^{\text{l}}(j)+\widetilde{\cos}_{k,r_{\max}}^{\text{r}}(j)\right)+b^{1/2}_{k}\widehat{c}_{k}\ ,\quad j\in V_{k}\ ,\quad k\in\mathds{N}\ .$
@@
$b_{k}$ $=2\left(1-b_{k}\right)^{1/2}b_{k}^{1/2}\widehat{c}_{k}\sum\limits_{j=-k}^{k}\left(\widetilde{\cos}_{k,r_{\min}}^{\text{l}}(j)+\widetilde{\cos}_{k,r_{\max}}^{\text{r}}(j)\right)+|V_{k}|b_{k}\widehat{c}_{k}^{2}\ .$ (20)
@@
Lastly, we prove (19). Note that there exists a constant $c>0$ such that
$\sum\limits_{j=-k}^{k}\big(\widetilde{\cos}_{k,r_{\min}}^{\text{l}}(j)+\widetilde{\cos}_{k,r_{\max}}^{\text{r}}(j)\big)\geq ck^{1/2}$ (21)
for all but finitely many $k\in\mathds{N}$. Let $k\in\mathds{N}$ be sufficiently large such that (21) and $c\widehat{\alpha}^{-1}k^{-1}\leq 3/4$ holds but otherwise arbitrary. Suppose that $b_{k}\geq c\widehat{\alpha}^{-1}k^{-1}$; then there is nothing left to show in regard to (19). Thus, suppose $b_{k}<c\widehat{\alpha}^{-1}k^{-1}\leq 3/4$. With (20) and (21) we conclude $b_{k}\geq cb_{k}^{1/2}\widehat{c}_{k}k^{1/2}$. Using (6) this then implies (19). ∎
@@
Lemma 6 (Bounds on first excited state energy).
For all but finitely many $k\in\mathds{N}$, one has
@@
The lower bound is a direct consequence of $\lambda_{1}(V_{k},\boldsymbol{\alpha})\geq\lambda_{1}(V_{k},\boldsymbol{0})=\lambda_{0}(V_{k},\infty\delta_{0})$. The upper bound follows from the operator inequality $\mathcal{L}_{k}+\sum_{j\in J}\alpha_{j}\delta_{j}\leq\mathcal{L}_{k}+\sum_{j\in J}\infty\delta_{j}$ in combination with $\lambda_{1}(V_{k},\sum_{j\in J}\infty\delta_{j})\leq\max\{\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0}),\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})\}$, which holds for all but finitely many $k\in\mathds{N}$. ∎
4. Main results I: General Case
@@
$\left(|V_{k}|^{3}\,\Gamma(V_{k},\boldsymbol{\alpha})\right)_{k\in\mathds{N}}$ (22)
and prove a conjecture regarding its limiting behavior as put forward in [KY25] (additional results will be provided in Section 5). By doing this and by considering potentials of compact support, we also generalize the main result of [KY25] – both – in terms of convergence speed and applicability to a broader class of potentials: more explicitly, we will show that
@@
for all $0\leq\eta<1$, as soon as the external potential is not the zero potential. Here, one should recall that $\lim_{k\rightarrow\infty}\left(|V_{k}|^{2}\cdot\Gamma(V_{k},\boldsymbol{\alpha})\right)=\pi^{2}$ for $\boldsymbol{\alpha}=\boldsymbol{0}$; compare with [KY25]. In other words, as soon as there is a non-vanishing external potential of compact support, the spectral gap converges strictly faster to zero than in the absence of any potential. This might be surprising at first, since the potential is supported on a smaller and smaller fraction of the configuration space in the limit of infinite volume. Consequently, this is a rather remarkable spectral effect.
We now state the first main result of this paper.
@@
    There exists a constant $C=C(\boldsymbol{\alpha})>0$ such that
    $|V_{k}|^{3}\,\Gamma(V_{k},\boldsymbol{\alpha})\leq C\quad\text{for all}\quad k\in\mathds{N}\ .$
@@
    If $\widehat{\alpha}=\sum_{j\in J}\alpha_{j}$ is sufficiently small or if $J=\{0\}$, there exists a constant $c=c(\boldsymbol{\alpha})>0$ such that
    $|V_{k}|^{3}\,\Gamma(V_{k},\boldsymbol{\alpha})\geq c\quad\text{for all}\quad k\in\mathds{N}\ .$
@@
$|V_{k}|^{3}\,\Gamma(V_{k},\boldsymbol{\alpha})\leq(const.)\widetilde{\alpha}^{-1}k^{-1}\Theta_{\max}|V_{k}|^{3}+\left(\Theta_{\max}-\Theta_{\min}\right)|V_{k}|^{3}\ .$ (23)
Now, taking into account that $\Theta_{\max}\sim k^{-2}$ and $\Theta_{\max}-\Theta_{\min}\lesssim k^{-3}$ completes the proof; recall here (5).
@@
$|V_{k}|^{3}\,\Gamma(V_{k},\boldsymbol{\alpha})\geq(const.)\widehat{\alpha}^{-1}\lambda_{0}(V_{k},\infty\delta_{0})k^{-1}|V_{k}|^{3}+\left(\lambda_{0}(V_{k},\infty\delta_{0})-\Theta_{\max}\right)|V_{k}|^{3}\ .$ (24)
@@
    Regarding the upper bound (23): Note that the first term is of the form $(const.)\widetilde{\alpha}^{-1}$ and consequently converges to zero when the potential strength converges to infinity in the sense that $\widetilde{\alpha}\to\infty$. However, the last term is independent of the potential strength and is non-zero if and only if $r_{\max}\neq-r_{\min}$. Therefore, the entire upper bound is of the form $(const.)\widetilde{\alpha}^{-1}$ and consequently converges to zero as $\widetilde{\alpha}$ converges to infinity if and only if $r_{\max}=-r_{\min}$.
@@
    Regarding the lower bound (24): The last term is negative whenever $J\neq\{0\}$ and it vanishes if and only if $J=\{0\}$; note that, trivially, $\Gamma(V_{k},\boldsymbol{\alpha})\geq 0$ for all $k\in\mathds{N}$. Therefore, this lower bound is non-trivial if and only if $\widehat{\alpha}$ is sufficiently small or if $J=\{0\}$. Furthermore, the first term of this bound is of the form $(const.)\widehat{\alpha}^{-1}$. Consequently, if and only if $J=\{0\}$, the entire lower bound is of the form $(const.)\widehat{\alpha}^{-1}$ and converges to zero as the strength of the potential tends to infinity in the sense that $\widehat{\alpha}\to\infty$. Note that the lower bound converges to $\infty$ as $\hat{\alpha}\to 0$. This refers to the case where we approach the free Laplacian, for which the spectral gap vanishes in lower order.
@@
    The reason we end up with a trivial lower bound when $J\neq\{0\}$ and $\widehat{\alpha}$ is too large is the lack of a sufficiently good lower bound for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ in terms of $\lambda_{0}(V_{k+r_{\min}},\infty\delta_{0})$ and/or $\lambda_{0}(V_{k-r_{\max}},\infty\delta_{0})$; compare with Lemma 6. Indeed, also from the physical point of view outlined in Remark 2, it is reasonable to expect that the lower bound for $\lambda_{1}(V_{k},\boldsymbol{\alpha})$ in Lemma 6 is suboptimal for strong external potentials. Nevertheless, a corresponding statement as in $(ii)$ of Theorem 7 is expected for all potential strengths.
Next, for the convenience of the reader, we summarize our findings from Theorem 7 in the special case $J=\{0\}$, that is, when the potential is localized on the zero vertex in the middle of the path graph. In addition, we provide upper and lower bounds for the limit that are explicit in the coupling strength. This special case was discussed in [KY25] and originally motivated this paper, and Corollary 9 proves the conjecture that was put forward in [KY25, Section 4].
@@
Consider the Hamiltonian $H_{k,\alpha}=\mathcal{L}_{k}+\alpha\delta_{0}$ with arbitrary $\alpha>0$ and associated spectral gap $\Gamma(V_{k},\alpha)$. Then there exist constants $c,C>0$ independent of $k$ and $\alpha$ such that for all but finitely many $k\in\mathds{N}$,
$\dfrac{c}{\alpha}\leq|V_{k}|^{3}\,\Gamma(V_{k},\alpha)\leq\dfrac{C}{\alpha}\ .$
@@
$|V_{k}|^{3}\,\Gamma(V_{k},\alpha)\leq(const.)\alpha^{-1}k^{-1}\lambda_{0}(V_{k},\infty\delta_{0})|V_{k}|^{3}\ .$
@@
$|V_{k}|^{3}\,\Gamma(V_{k},\alpha)\geq(const.)\alpha^{-1}k^{-1}\lambda_{0}(V_{k},\infty\delta_{0})|V_{k}|^{3}$